# Tessellated phase diagram and extended criticality in driven quasicrystals and quantum Hall matter

### v2

@@ II.3 Floquet engineering of multifractality without fine-tuning
In this section, we will show that these control capabilities afforded by the DDAAH model (2) allow preparation of an extended critical phase of matter hosting multifractal wavefunctions without fine-tuning [8, 10, 11]. The possibility of an extended critical phase is of interest from both theoretical and experimental perspectives: usually, multifractal wave functions only appear when a system is finely tuned to a phase boundary or mobility edge, which hinders experimental study.
Earlier work on irradiated graphene showed that elliptically polarized irradiation introduces NNN hopping in the effective Hamiltonian through absorption and emission of virtual photons [46]. Moreover, when these NNN hoppings are present in the static 2D Harper-Hofstadter model with an incommensurate flux, they are known to introduce critical phases [8, 55]. As we will explicitly demonstrate below, it is therefore possible to engineer critical phases by applying elliptical irradiation to the Harper-Hofstadter Hamiltonian in the 2D space (see Fig. 5(a)).
Equipped with this observation, we apply the elliptical modulation to the 1D physical system. For simplicity, here we consider only the special case of circularly polarized modulation ($\delta\theta=\pm\pi/2$ and $K_{0}=\varphi_{0}$), and defer the discussion of general polarization to Appendix B. Then $\hat{H}^{(1)}\neq 0$ and the effective Hamiltonian is $\hat{H}_{\mathrm{eff}}\approx\hat{H}^{(0)}+\hat{H}^{(1)}$, where
@@
An important question in this context concerns the role of higher-order terms ($\hat{H}^{(n\geq 2)}$) in the high-frequency expansion. These terms, though small in the high-frequency regime, can potentially remove the multifractality of certain Floquet eigenstates in the spectrum, partly because these terms can break the delicate generalized self-duality of the extended AAH model [57, 55]. To examine their impact, we numerically calculate the fractal dimension $D_{2}$ of the eigenstates of the Floquet block Hamiltonian instead of the effective Hamiltonian from the high-frequency expansion, thereby including every term in the high-frequency expansion in principle. The results, which appear in Fig. 6(b), show that although certain Floquet eigenstates cease to be multifractal when the full Floquet Hamiltonian is considered, a finite fraction of Floquet eigenstates still exhibit multifractal characteristics with $D_{2}\neq 0,1$. We further corroborate this result by numerical diagonalization of the Hamiltonian $\hat{H}^{(0)}+\hat{H}^{(1)}+\hat{H}^{(2)}$, which is discussed in the Appendix C.
We thus establish the existence of multifractality without fine-tuning in the DDAAH model when the combined modulations are tuned to correspond to circularly polarized light in the 2D space, and we further generalize this result to elliptical modulation in the Appendix B. While previously proposed methods for generating critical phases depend upon mixing localized and extended eigenstates of the undriven system using low-frequency periodic driving [25] or unbounded quasiperiodic potentials [24], our model contains neither of those elements, and is inherently in the high-frequency regime, which may alleviate interband heating in experimental platforms based on ultracold atoms [22]. The effective Hamiltonian up to $\hat{H}^{(1)}$ corresponds to a model that had eluded direct experimental realization until very recently with programmable superconducting qubits [29] and a constrained system size. We expect that the approach to achieving an extended fractal phase described in this section may significantly simplify experimental realization, as further discussed in section IV.
@@
III From driven Harper-Hofstadter to the generalized Su-Schrieffer-Heeger model
It is known that the 1D RM model and SSH models can be projected from the 2D Harper-Hofstadter model with NNN hoppings [60, 61, 62]. Consider the case where the 2D Harper-Hofstadter model with NNN hoppings has a magnetic flux $\beta=1/q$ (where $q$ is a positive integer) through one plaquette. The projected 1D model is a superlattice with $q$ sites per unit cell. In the projected space, both the on-site potential and the nearest-neighbor tunneling are sinusoidally modulated with the same spatial period $q$. The projected model then describes a RM model with $q$ sites per unit cell. When the on-site potential vanishes, the model becomes the SSH with $q$ sites per unit cell [60].
$\mathcal{B}=\sum_{j\in\{\mathrm{L},\mathrm{R}\}}|\psi_{j}^{(i)}|^{2}$ (9)
where L and R represent sites at left and the right edge, respectively.
$+\mathrm{h.c.}.$ (10)
IV Experimental realization
@@
Figure 9: Time evolution of the rms width for the DDAAH model (2) at different locations in the phase diagram. Parameters, the polarization of the modulations, the fitted values of the exponent $\gamma$ and the corresponding phases are shown at the top panel. The time-evolution is shown at the bottom panel, where dotted black lines represent the fit for the exponent $\gamma$. The system is driven for $2000$ cycles at $\hbar\omega=15J$ with $\varphi_{0}=K_{0}$ and data in each case are averaged over 10 randomly selected values of $\kappa$. The initial state is a single-site excitation $\psi(t)=\delta_{j,0}$ and the Hamiltonian (2) is numerically integrated with the second order split-operator method with a time step of $dt\approx 1.26\times 10^{-5}\hbar/J$.
@@
Circularly polarized modulation can be most easily achieved by shaking the primary lattice while keeping the secondary lattice static ($x_{S}(t)=0$) in the lab frame by introducing a frequency difference $f_{P}(t)=f_{0,P}\sin(\omega t)$ between the two counter-propagating lattice beams that form the primary lattice. In the co-moving frame, the dipolar and phasonic modulations are then given by [33]
@@
We demonstrate this usage of $\gamma$ by computing the time-evolution of the wave functions at exemplary points in the phase diagram, with the Hamiltonian in the length gauge (Eq. 2). The results in Fig. 9 show the expected ballistic expansion in the delocalized phase ($\gamma\approx 1$), the suppression of transport at long times in the localized phase ($\gamma\approx 0$), and the critical expansion with $\gamma\approx 0.4067$ when RCP modulation is applied. In contrast, when the modulation is linearly polarized, the same parameters no longer leads to the critical dynamic but a localized one.
V Summary and Outlook
In this work, we introduced combined dipolar and phasonic modulations to the AAH model, and outlined a mapping to a 2D integer quantum Hall system illuminated by light of variable polarization. The combined modulations offer a remarkable degree of control over localization and topological properties. Phenomena that can thereby be physically realized include tessellated phase diagrams with interlaced localized and delocalized phases, effective time-reversal dynamics, Floquet-engineered multifractality without fine-tuning, and topological edge states. In describing these phenomena we identified a connection between the elliptically driven Harper-Hofstatder model and generalized Rice-Mele and Su-Shrieffer-Heeger models.
These results pave the way for future work, in part because the simplicity of our model reduces experimental challenges. Here we outline some future opportunities. A straightforward extension is to use low-dimensional quasiperiodic systems to study the dynamics of corresponding strongly-driven integer quantum Hall systems in higher dimensions [74, 75, 76, 13, 42]. For example, one can generalize our results to study 4D driven quantum Hall systems with 2D doubly-driven bichromatic lattices [13, 42]. Extensions to a driven Hofstadter model with a non-abelian gauge field are also possible by considering spinful fermions in the DDAAH model [77]. Our results may have implications for driven strained Moiré superlattices in a magnetic field, where the anisotropic conductivity can be switched by irradiation (Fig. 3 and [78]). Finally, the inclusion of interparticle interactions poses a challenging and fundamental question: can one expect a Floquet many-body critical phase? In some static systems, critical phases have been shown numerically to retain their multifractal characteristics in the presence of interactions [24, 30], thus appearing as an additional possibility alongside thermalization and many-body localization. Probing the stability of the Floquet critical phase in the presence of interactions is thus a natural target for future investigation [79, 80].
@@ Appendix B High-frequency expansion of modulations with a generic polarization
The second difficulty comes with the computational cost associated with the system-size scaling. In the numerical diagonalization, the number of Floquet-Brillouin zones is $(2m_{\mathrm{cutoff}}+1)$, and the block matrix has a dimension of $(2m_{\mathrm{cutoff}}+1)L$ where $L$ is the system size. Consequently, the multifractal analysis with the Floquet block Hamiltonian quickly becomes numerically expensive as the system size $L$ is increased, even with a relatively small $m_{\mathrm{cutoff}}$.
@@
Here we compare and the contrast these two approaches by numerical calculation of the fractal dimension in the same parameter range. For the truncated high-frequency expansion we take $n_{\mathrm{cutoff}}=2$. The result is shown in Fig. 10. Values of $D_{2}$ are still non-trivial in a wide range of parameters and for a non-vanishing fraction of eigenstates, and there are very few differences between these two calculations. This further supports the existence of the critical phase in our model.

### v3

@@ II.3 Floquet engineering of multifractality without fine-tuning
In this section, we will show that the control capabilities afforded by tuning the drive polarization in the DDAAH model (2) allow preparation of an extended critical phase of matter hosting multifractal wavefunctions without fine-tuning [8, 10, 11]. The possibility of an extended critical phase is of interest from both theoretical and experimental perspectives; usually, multifractal wave functions only appear when a system is finely tuned to a phase boundary or mobility edge, which hinders experimental study.
Previous work on irradiated graphene has shown that elliptically polarized irradiation introduces NNN hopping in the effective Hamiltonian through absorption and emission of virtual photons [48]. Moreover, when these NNN hoppings are present in the static 2D Harper-Hofstadter model with an incommensurate flux, they are known to introduce critical phases [8, 57]. As we will explicitly demonstrate below, it is therefore possible to engineer critical phases by applying elliptical irradiation to the Harper-Hofstadter Hamiltonian in the 2D space (see Fig. 5(a)).
Motivated by this observation, we apply elliptical modulation to the 1D physical system. For simplicity, here we consider only the special case of circularly polarized modulation ($\delta\theta=\pm\pi/2$ and $K_{0}=\varphi_{0}$), and defer the discussion of general polarization to Appendix B. Then $\hat{H}^{(1)}\neq 0$ and the effective Hamiltonian is $\hat{H}_{\mathrm{eff}}\approx\hat{H}^{(0)}+\hat{H}^{(1)}$, where
@@
An important question in this context concerns the role of higher-order terms ($\hat{H}^{(n\geq 2)}$) in the high-frequency expansion. These terms, though small in the high-frequency regime, can potentially remove the multifractality of certain Floquet eigenstates in the spectrum, partly because these terms can break the delicate generalized self-duality of the extended AAH model [59, 57]. To examine their impact, we numerically calculate the fractal dimension $D_{2}$ of the eigenstates of the Floquet block Hamiltonian instead of the effective Hamiltonian from the high-frequency expansion, thereby including every term in the high-frequency expansion in principle. The results, which appear in Fig. 6(b), show that although certain Floquet eigenstates cease to be multifractal when the full Floquet Hamiltonian is considered, a finite fraction of Floquet eigenstates still exhibit multifractal characteristics with $D_{2}\neq 0,1$. We further corroborate this result by numerical diagonalization of the Hamiltonian $\hat{H}^{(0)}+\hat{H}^{(1)}+\hat{H}^{(2)}$, which is discussed in the Appendix C. We note, however, that because $\hat{H}^{(1)}$ is still the dominant contribution, quantities like the averaged IPR still serve as a meaningful metric despite the small mixing of the spectrum.
We thus establish the existence of multifractality without fine-tuning in the DDAAH model when the combined modulations are tuned to correspond to circularly polarized light in the 2D space, and we further generalize this result to elliptical modulation in the Appendix B. While previously proposed methods for generating critical phases depend upon mixing localized and extended eigenstates of the undriven system using low-frequency periodic driving [25] or unbounded quasiperiodic potentials [24], our model contains neither of those elements, and is inherently in the high-frequency regime, which should significantly alleviate the problem of interband heating in experimental platforms based on ultracold atoms [22]. The effective Hamiltonian up to $\hat{H}^{(1)}$ corresponds to a model that had eluded direct experimental realization until very recently, with programmable superconducting qubits [29] and a constrained system size. We expect that the approach to achieving an extended fractal phase described in this section may significantly simplify experimental realization, as further discussed in section III.
@@
III Experimental realization
@@
Figure 7: Time evolution of the rms width (Eq. 10) at different locations on the phase diagram in the space of drive polarization. The inset shows the phase diagram under circularly polarized modulation with markers at the locations where the numerical simulations are conducted. The top panel shows the parameters, polarization, the expansion exponent $\gamma$ and the underlying phases. The bottom panel shows the time evolution of the width $\sigma(t)$ from which the exponents $\gamma$ are obtained by fitting, shown by the black lines. The system is driven for $2000$ cycles at $\hbar\omega=15J$ with $\varphi_{0}=K_{0}$ and data in each case are averaged over 10 randomly selected values of $\kappa$. The initial state is a single-site excitation $\psi(t)=\delta_{j,0}$ and the Hamiltonian (2) is numerically integrated with the second order split-operator method with a time step of $dt\approx 1.26\times 10^{-5}\hbar/J$.
@@
Circularly polarized modulation can be achieved most easily by shaking the primary lattice while keeping the secondary lattice static ($x_{S}(t)=0$) in the lab frame by introducing a frequency difference $f_{P}(t)=f_{0,P}\sin(\omega t)$ between the two counterpropagating lattice beams that form the primary lattice. In the co-moving frame, the dipolar and phasonic modulations are then given by [33]
@@
We theoretically demonstrate the potential use of this experimental signature by computing the time evolution of the wavefunction at exemplary points in the phase diagram, using the Hamiltonian in the length gauge (Eq. 2). The results in Fig. 7 show the expected ballistic expansion in the delocalized phase ($\gamma\approx 1$), the suppression of transport at long times in the localized phase ($\gamma\approx 0$), and the critical expansion with $\gamma\approx 0.4067$ when circularly polarized modulation is applied. In contrast, when the modulation is linearly polarized, the same values of $\Delta$ and $K_{0}$ no longer lead to critical dynamics but rather to localization; this provides a particularly clean demonstration of control of the state of matter using drive polarization.
IV Summary and Outlook
In this work, we introduced combined dipolar and phasonic modulations to the AAH model, and outlined a mapping to a 2D integer quantum Hall system illuminated by light of variable polarization. The combined modulations offer a remarkable degree of control over localization. Phenomena that can thereby be physically realized include tessellated phase diagrams with interlaced localized and delocalized phases, effective time-reversal dynamics, Floquet-engineered multifractality without fine-tuning, and (as described in appendix D) topological edge states.
These results pave the way for future work, in part because the simplicity of our model lowers the barrier to experimental realization. Here we outline some future opportunities.
A straightforward extension is to use low-dimensional quasiperiodic systems to study the dynamics of corresponding strongly-driven integer quantum Hall systems in higher dimensions [69, 70, 71, 13, 44]. For example, one can generalize our results to study 4D driven quantum Hall systems with 2D doubly-driven bichromatic lattices [13, 44]. Extensions to a driven Hofstadter model with a non-abelian gauge field are also possible by considering spinful fermions in the DDAAH model [72]. Our results may have implications for driven strained Moiré superlattices in a magnetic field, where the anisotropic conductivity can be switched by irradiation (Fig. 3 and [73, 55]). While we focus on off-resonant driving in this work, resonant driving could allow us to study quantum geometry of a higher-dimensional system using its lower-dimensional counterpart. Possible targets of study include circular dichroism in one dimension [74] and quantum metrics [75]. Finally, the inclusion of interparticle interactions poses a challenging and fundamental question: can one expect a Floquet many-body critical phase? In some static systems, critical phases have been shown numerically to retain their multifractal characteristics in the presence of interactions [24, 30], thus appearing as an additional possibility alongside thermalization and many-body localization. Probing the stability of the Floquet critical phase in the presence of interactions is thus a natural target for future investigation [76, 77].
@@ Appendix B High-frequency expansion of modulations with a generic polarization
An associated practical challenge is the computational cost associated with the system-size scaling. In the numerical diagonalization, the number of Floquet-Brillouin zones is $(2m_{\mathrm{cutoff}}+1)$, and the block matrix has a dimension of $(2m_{\mathrm{cutoff}}+1)L$ where $L$ is the system size. Consequently, the multifractal analysis with the Floquet block Hamiltonian quickly becomes numerically expensive as the system size $L$ is increased, even for a relatively small $m_{\mathrm{cutoff}}$.
@@
Here we compare and the contrast these two approaches by numerical calculation of the fractal dimension in the same parameter range. For the truncated high-frequency expansion we take $n_{\mathrm{cutoff}}=2$. The result is shown in Fig. 8. Values of $D_{2}$ are still non-trivial in a wide range of parameters and for a non-vanishing fraction of eigenstates, and we observe only minor differences between the results of these two calculations. This further supports the existence of the critical phase in our model.
Appendix D From driven Harper-Hofstadter to the generalized Su-Schrieffer-Heeger model
It is known that the 1D RM model and SSH models can be projected from the 2D Harper-Hofstadter model with NNN hoppings [81, 82, 83]. Consider the case where the 2D Harper-Hofstadter model with NNN hoppings has a magnetic flux $\beta=1/q$ (where $q$ is a positive integer) through one plaquette. The projected 1D model is a superlattice with $q$ sites per unit cell. In the projected space, both the on-site potential and the nearest-neighbor tunneling are sinusoidally modulated with the same spatial period $q$. The projected model then describes a RM model with $q$ sites per unit cell. When the on-site potential vanishes, this becomes the SSH model with $q$ sites per unit cell [81].
$\mathcal{B}=\sum_{j\in\{\mathrm{L},\mathrm{R}\}}|\psi_{j}^{(i)}|^{2}$ (13)
where L and R represent sites at left and the right edge, respectively.
$+\mathrm{h.c.}.$ (14)